# Parameter Efficient Continual Learning with Dynamic Low-Rank Adaptation

### v3

@@ B.5 Hyperparameter Tuning
B.6 Comparison against CL approaches in standard benchmarks.
@@
PNNs [rusu2016progressive] NeurIPS 2016 - 74.01 $\pm$1.11 - 67.84 $\pm$0.29
PackNet [mallya2018packnet] CVPR 2018 - 72.39 $\pm$0.37 - 60.46 $\pm$1.22
NISPA [gurbuz2022nispa] NeurIPS 2022 - 65.36 $\pm$2.19 - 59.56 $\pm$0.32
SparCL-EWC [wang2022sparcl] NeurIPS 2022 - 59.53 $\pm$0.25 - 59.56 $\pm$0.32
@@
Co2L [cha2021co2l] ICCV 2021 31.90 $\pm$0.38 55.02 $\pm$0.36 13.88 $\pm$0.40 42.37 $\pm$0.74
CLS-ER [arani2022learning] ICLR 2022 43.80 $\pm$1.89 73.49 $\pm$1.04 23.47 $\pm$0.80 49.60 $\pm$0.72
OCDNet [li2022learning] IJCAI 2022 44.29 $\pm$0.49 73.53 $\pm$0.24 17.60 $\pm$0.97 56.19 $\pm$1.31
TAMiL [bhat2022task] ICLR 2022 41.43 $\pm$0.75 71.39 $\pm$0.17 20.46 $\pm$0.40 55.44 $\pm$0.52
IMEX-Reg [bhat_imex-reg_2024] TMLR 2024 48.54 $\pm$0.23 75.61 $\pm$0.73 24.15 $\pm$0.78 62.91 $\pm$0.54
TriRE [TriRE2023] NeurIPS 2023 43.91 $\pm$0.18 71.66 $\pm$ 0.44 20.14 $\pm$0.19 55.95 $\pm$0.78
AGILE [bhat2024mitigating] CoLLAs 2024 45.73 $\pm$0.15 74.37 $\pm$0.34 20.19 $\pm$1.65 53.47 $\pm$1.60
SARL [sarfrazsemantic] ICLR 2025 48.96 $\pm$ 0.53 78.91 $\pm$ 0.15 28.95 $\pm$ 1.13 71.56 $\pm$0.18
@@
B.7 Performance in longer task sequences
@@
B.8 Growth in model size
Dataset Backbone Method 5 Tasks 10 Tasks 20 Tasks
Seq-CIFAR100 R18 Complete Param. Isolation 55.89 111.73 223.42
PEARL 12.56 13.27 13.55
BSC Complete Param. Isolation 6.32 12.60 25.14
PEARL 1.63 1.89 2.30
ImageNet-R ViT Complete Param. Isolation 452.86 905.72 1,811.44
PEARL 92.46 93.27 95.35
B.9 Dynamic rank allocation across layers in ViT
PEARL introduces a PEFT CL framework that incorporates dynamic rank selection for LoRA components during training. PEARL leverages reference task weights and adaptively determines the rank of task-specific, layer-specific LoRA components based on the proximity of the current task to the reference task weights in the parameter space. Figure 1 illustrates the dynamic rank allocation across layers for convolutional variants for the final task in Seq-CIFAR100 (5T). From left to right, Figure 1 presents a detailed overview of the layer-wise dynamic thresholds established based on task proximity, along with the corresponding selected rank and actual rank for our convolutional variants. Similarly, Figure 4 presents the dynamic rank allocation across layers for the final task in ImageNet-R (5T) for PEARL (ViT). In the ViT model, earlier layers are primarily responsible for extracting low-level image features. These features lay the foundation for subsequent high-level feature extraction and semantic understanding in later layers. As these features relatively less complex and can be re-used across tasks, PEARL puts less emphasis on them through lower rank allocation. On the other hand, intermediate layers handle more abstract and complex features, such as parts of objects and entire objects, necessitating higher rank allocation. Therefore, PEARL puts a major emphasis on intermediate layers with higher rank allocation. The top layers of the ViT model are designed to capture high-level semantic features and global contextual information. By later layers, the model has likely already captured the most significant semantic features relevant to the task, and additional layers contribute marginally smaller improvements to the final output quality. Therefore, PEARL allocates a lower threshold and a corresponding lower rank in later layers. In line with the aforementioned findings in zeng2024peeling, PEARL displays a bell-curve shaped dynamic rank allocation across layers thereby optimizing resource allocation and maximizing information reuse.
@@ B.9 Dynamic rank allocation across layers in ViT
B.10 Effect of random class ordering
The Table 12 displays the Class-IL performance on ImageNet-R (10 tasks) under five different class orders, which differ from those presented in Table 2. In this setting, we randomize class assignments across tasks by setting different random seeds, allowing us to assess the impact of class ordering on PEARL. Although the final Class-IL accuracy ($A_{T}$%) is slightly lower than the previously reported result of $88.76\pm 0.69$, the results remain consistent and robust across these new class orders, highlighting the method’s stability. Furthermore, both incremental and final accuracies show stability across varying class orderings, reinforcing the reliability of the approach. Even across all different class orderings, PEARL (ViT) consistently outperforms the considered baselines in Table 2 by a huge margin. In conclusion, the findings indicate that PEARL maintains robust performance and generalizability across varying class orders, affirming its reliability as an effective framework for CL, even in randomized conditions.
Class Order # Params (Millions) $_{T}A¯\bar{A}$%
1 92.78 86.63 91.24
2 93.46 86.76 92.24
3 93.41 86.50 91.54
4 93.26 86.73 91.69
5 92.83 86.86 91.77
Average 93.14 86.70 91.70
B.11 Effect of pre-training on PEARL
Method Seq-CIFAR100 (5T) Seq-TinyImageNet (10T)
$_{T}A¯\bar{A}$% $_{T}A¯\bar{A}$%
B.12 Additional comparison with adapter-based methods
@@
Table 14: Following [10658306], we compare additional adapter-based methods in the Class-IL setting on ImageNet-R (10T) using a ViT-B/16 backbone pre-trained on ImageNet-21K.
@@
Adam-ssf [10.1007/s11263-024-02218-0] 66.61 $\pm$ 0.09 74.36 $\pm$ 1.00 0.20
LAE [gao2023lae] 72.29 $\pm$ 0.14 77.99 $\pm$ 0.46 0.19
ADA [10.5555/3600270.3601042] 73.76 $\pm$ 0.27 79.57 $\pm$ 0.84 1.19
SSIAT [10658306] 79.38 $\pm$ 0.59 83.63 $\pm$ 0.43 1.19
@@
B.13 Additional comparison with prompt-based methods
Table 15 provides an extensive comparison of prompt-based methods in the continual learning (CL) setting on ImageNet-R. Following HiDe-Prompt [wang2023hide], we evaluate our method using the final forgetting measure (FFM) and cumulative forgetting measure (CFM), defined as follows. Let $A_{i,t}$ denote the accuracy on the $i$-th task after learning the $t$-th task.
@@ B.13 Additional comparison with prompt-based methods
Table 15: Following [wang2023hide], we compare additional prompt-based methods in the Class-IL setting on ImageNet-R using a ViT-B/16 backbone pre-trained on ImageNet-21K. Bold indicates the best result and underline indicates the second-best result.
@@
L2P [wang2022learning] $3.94\pm 0.16$ $3.55\pm 0.20$ $5.01\pm 0.40$ $4.41\pm 0.43$ $10.76\pm 0.45$ $7.88\pm 0.17$
DualPrompt [DBLP:conf/eccv/0002ZESZLRSPDP22] $3.32\pm 0.16$ $2.78\pm 0.25$ $4.61\pm 0.07$ $3.70\pm 0.18$ $7.30\pm 0.18$ $5.16\pm 0.34$
CODA-P [smith2023coda] $8.89\pm 0.65$ $7.65\pm 0.98$ $7.94\pm 0.08$ $6.72\pm 0.79$ $8.23\pm 0.86$ $6.95\pm 0.70$
LGCL [pub13491] $3.04\pm 0.36$ $2.50\pm 0.38$ $4.75\pm 0.33$ $3.38\pm 0.58$ $7.35\pm 0.65$ $5.05\pm 0.32$
@@
PGP [qiao2024prompt] $3.36\pm 0.23$ $2.85\pm 0.25$ $4.53\pm 0.40$ $3.63\pm 0.35$ $7.17\pm 0.21$ $5.09\pm 0.25$
EvoPrompt [Kurniawan_Song_Ma_He_Gong_Qi_Wei_2024] $\mathbf{1.79\pm 0.31}$ $\underline{1.41\pm 0.32}$ $4.22\pm 0.42$ $3.59\pm 0.52$ $6.72\pm 0.90$ $5.67\pm 0.26$
CPrompt [gao2024consistent] $6.00\pm 0.00$ $5.49\pm 0.00$ $6.10\pm 0.75$ $5.60\pm 1.35$ $98\pm 0.24 5.5$$54\pm 0.48$
ConvPrompt [roy_2024_CVPR] $3.42\pm 0.05$ $2.36\pm 0.16$ $4.17\pm 0.04$ $3.11\pm 0.17$ $\mathbf{4.87\pm 0.57}$ $3.57\pm 0.25$
@@
B.14 Inference time comparison in Class-IL
@@
L2P [wang2022learning] 9.53
DualPrompt [wang2022dualprompt] 9.51
CODA-P [smith2023coda] 9.47
Adam-ssf [10.1007/s11263-024-02218-0] 10.03
SEMA [SEMA] 7.39
@@
B.15 Task-wise resource allocation
@@
Appendix C Experimental setup
C.1 Backbones
To demonstrate the versatility of PEARL across different architectures, we also explore three distinct backbones: (i) Convolutional networks such as ResNet-18 and a ResNet-18-like architecture incorporating Blueprint Separable Convolutions (BSC) [li2022blueprint], and (ii) the Vision Transformer (ViT) Base 16 pre-trained on ImageNet-21k.
With regard to the PEARL (R18), the backbone is same as the one used in Mammoth CL benchmark [buzzega2020dark]. In line with previous works, the backbone is randomly initialized and trained for CL tasks. However, we introduce PyTorch ModuleLists all throughout to make it easier to add / remove convolutional and BatchNorm layers at will. The ResNet-18 backbone consists of 4 layers with 2 blocks each. Each block in turn consists of two 3x3 convolutional layers with a BatchNorm and ReLU activation in between and after. ResNet-18 also entails shortcut connections in the form of a 3x3 convolutional layer followed by a BatchNorm layer and ReLU activation. Within PEARL (R18), we reserve separate classification and BatchNorm layers for each task. As the backbone is not pre-trained, we treat the first task as the reference task and conduct dynamic allocation from the second task onward. LoRA components are allocated for every convolutional layer. Their rank is dynamically decided after SVD decomposition of respective task vectors. Even after dynamic allocation and further training, LoRA components, BatchNorm layers and a corresponding classification layer are kept separate for each task.
li2022blueprint introduced blueprint separable convolutions (BSC) to optimize and reduce redundancy in convolutional operations. Essentially, BSC entails a 1x1 operation (point-wise filter) followed by a depth-wise convolution layer. This separability reduces the memory and computational footprint resulting far more efficient model. Such a model is well suited for parameter isolation CL approach as it reduces overall model size when dealing with longer task sequences. PEARL (BSC) backbone consists of 3 layers with 2 blocks each. Within each block, the organization of layers is same as that of ResNet-18 except that traditional convolutional layers are replaced by BSC layers. As the BSC backbone is a custom, small network, pre-trained model is not readily available. Therefore, we treat the first task as the reference task and conduct dynamic allocation from the second task onward. LoRA components are allocated for every point-wise filters as they have the majority of the footprint in terms of number of parameters. Since depth-wise filters have relatively lower footprint, we allocate LoRA components and select the rank of the task vector after decomposition. Same as in ResNet-18, even after dynamic allocation and further training, LoRA components, BatchNorm layers and a corresponding classification layer are kept separate for each task.
@@ C.1 Backbones
C.2 Datasets and settings
We evaluate PEARL across two distinct scenarios of CL: Class-IL and Task-IL. In both these settings, the model encounters a new set of classes in each task and must learn to distinguish all classes encountered thus far after each task. In practice, we split CIFAR-100 [krizhevsky2009learning], TinyImageNet [le2015tiny], and ImageNet-R [hendrycks2021many] into different tasks depending on the number of tasks as in Table 17.
@@ C.2 Datasets and settings
C.3 Training details
For experiments concerning PEARL (ViT) on the ImageNet-R dataset, we use the Adam optimizer [kingma2015adam] with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The batch size is set to 128. The ViT model is fine-tuned on the current task with a learning rate of $5\times 10^{-3}$ for 25 epochs ($E_{1}$). Following this, the LoRA weights are trained separately for an additional 25 epochs ($E_{2}$) using a learning rate of $5\times 10^{-4}$. We observed that a cosine-decaying learning rate schedule outperforms a constant learning rate. Regarding the LoRA parameterization, the rank is determined dynamically based on the criteria outlined in our approach, and the LoRA alpha parameter is set to twice the rank.
With regard to PEARL convolutional variants, we use a batch size of 32 for all experiments. Furthermore, we fine-tune the reference task weights for 15 epochs ($E_{1}$) with a best learning rate found in Section B.5. Following this, the LoRA components are initialized with dynamic ranks and are fine-tuned further for 50 epochs ($E_{2}$).

### v4

@@ C.3 Hyperparameter Tuning
C.4 Comparison against CL approaches in standard benchmarks.
@@
PNNs (Rusu et al., 2016) NeurIPS 2016 - 74.01 $\pm$1.11 - 67.84 $\pm$0.29
PackNet (Mallya & Lazebnik, 2018) CVPR 2018 - 72.39 $\pm$0.37 - 60.46 $\pm$1.22
NISPA (Gurbuz & Dovrolis, 2022) NeurIPS 2022 - 65.36 $\pm$2.19 - 59.56 $\pm$0.32
SparCL-EWC (Wang et al., 2022b) NeurIPS 2022 - 59.53 $\pm$0.25 - 59.56 $\pm$0.32
@@
Co2L (Cha et al., 2021) ICCV 2021 31.90 $\pm$0.38 55.02 $\pm$0.36 13.88 $\pm$0.40 42.37 $\pm$0.74
OCDNet (Li et al., 2022a) IJCAI 2022 44.29 $\pm$0.49 73.53 $\pm$0.24 17.60 $\pm$0.97 56.19 $\pm$1.31
TAMiL (Bhat et al., 2023) ICLR 2022 41.43 $\pm$0.75 71.39 $\pm$0.17 20.46 $\pm$0.40 55.44 $\pm$0.52
TriRE (Vijayan et al., 2023a) NeurIPS 2023 43.91 $\pm$0.18 71.66 $\pm$ 0.44 20.14 $\pm$0.19 55.95 $\pm$0.78
AGILE (Bhat et al., 2024) CoLLAs 2024 45.73 $\pm$0.15 74.37 $\pm$0.34 20.19 $\pm$1.65 53.47 $\pm$1.60
@@
C.5 Performance in longer task sequences
@@
C.6 Dynamic rank allocation across layers in ViT
PEARL introduces a PEFT CL framework that incorporates dynamic rank selection for LoRA components during training. PEARL leverages reference task weights and adaptively determines the rank of task-specific, layer-specific LoRA components based on the proximity of the current task to the reference task weights in the parameter space. Figure 2 illustrates the dynamic rank allocation across layers for convolutional variants for the final task in Seq-CIFAR100 (5T). From left to right, Figure 2 presents a detailed overview of the layer-wise dynamic thresholds established based on task proximity, along with the corresponding selected rank and actual rank for our convolutional variants. Similarly, Figure 4 presents the dynamic rank allocation across layers for the final task in ImageNet-R (5T) for PEARL (ViT). In the ViT model, earlier layers are primarily responsible for extracting low-level image features. These features lay the foundation for subsequent high-level feature extraction and semantic understanding in later layers. As these features relatively less complex and can be re-used across tasks, PEARL puts less emphasis on them through lower rank allocation. On the other hand, intermediate layers handle more abstract and complex features, such as parts of objects and entire objects, necessitating higher rank allocation. Therefore, PEARL puts a major emphasis on intermediate layers with higher rank allocation. The top layers of the ViT model are designed to capture high-level semantic features and global contextual information. By later layers, the model has likely already captured the most significant semantic features relevant to the task, and additional layers contribute marginally smaller improvements to the final output quality. Therefore, PEARL allocates a lower threshold and a corresponding lower rank in later layers. In line with the aforementioned findings in Zeng et al. (2024), PEARL displays a bell-curve shaped dynamic rank allocation across layers thereby optimizing resource allocation and maximizing information reuse.
@@ C.6 Dynamic rank allocation across layers in ViT
C.7 Additional comparison with adapter-based methods
@@
Table 14: Following Tan et al. (2024), we compare additional adapter-based methods in the Class-IL setting on ImageNet-R (10T) using a ViT-B/16 backbone pre-trained on ImageNet-21K.
@@
Adam-ssf (Zhou et al., 2024a) 66.61 $\pm$ 0.09 74.36 $\pm$ 1.00 0.20
LAE (Gao et al., 2023a) 72.29 $\pm$ 0.14 77.99 $\pm$ 0.46 0.19
ADA (Ermis et al., 2022) 73.76 $\pm$ 0.27 79.57 $\pm$ 0.84 1.19
SSIAT (Tan et al., 2024) 79.38 $\pm$ 0.59 83.63 $\pm$ 0.43 1.19
@@
C.8 Additional comparison with prompt-based methods
Table 15 provides an extensive comparison of prompt-based methods in the continual learning (CL) setting on ImageNet-R. Following HiDe-Prompt (Wang et al., 2023a), we evaluate our method using the final forgetting measure (FFM) and cumulative forgetting measure (CFM), defined as follows. Let $A_{i,t}$ denote the accuracy on the $i$-th task after learning the $t$-th task.
@@ C.8 Additional comparison with prompt-based methods
Table 15: Following Wang et al. (2023a), we compare additional prompt-based methods in the Class-IL setting on ImageNet-R using a ViT-B/16 backbone pre-trained on ImageNet-21K. Bold indicates the best result and underline indicates the second-best result.
@@
L2P (Wang et al., 2022e) $3.94\pm 0.16$ $3.55\pm 0.20$ $5.01\pm 0.40$ $4.41\pm 0.43$ $10.76\pm 0.45$ $7.88\pm 0.17$
DualPrompt (Wang et al., 2022d) $3.32\pm 0.16$ $2.78\pm 0.25$ $4.61\pm 0.07$ $3.70\pm 0.18$ $7.30\pm 0.18$ $5.16\pm 0.34$
CODA-P (Smith et al., 2023b) $8.89\pm 0.65$ $7.65\pm 0.98$ $7.94\pm 0.08$ $6.72\pm 0.79$ $8.23\pm 0.86$ $6.95\pm 0.70$
LGCL (Khan et al., 2023) $3.04\pm 0.36$ $2.50\pm 0.38$ $4.75\pm 0.33$ $3.38\pm 0.58$ $7.35\pm 0.65$ $5.05\pm 0.32$
@@
PGP Qiao et al. (2024) $3.36\pm 0.23$ $2.85\pm 0.25$ $4.53\pm 0.40$ $3.63\pm 0.35$ $7.17\pm 0.21$ $5.09\pm 0.25$
EvoPrompt (Kurniawan et al., 2024) $\mathbf{1.79\pm 0.31}$ $\underline{1.41\pm 0.32}$ $4.22\pm 0.42$ $3.59\pm 0.52$ $6.72\pm 0.90$ $5.67\pm 0.26$
CPrompt (Gao et al., 2024) $6.00\pm 0.00$ $5.49\pm 0.00$ $6.10\pm 0.75$ $5.60\pm 1.35$ $98\pm 0.24 5.5$$54\pm 0.48$
ConvPrompt (Roy et al., 2024) $3.42\pm 0.05$ $2.36\pm 0.16$ $4.17\pm 0.04$ $3.11\pm 0.17$ $\mathbf{4.87\pm 0.57}$ $3.57\pm 0.25$
@@
C.9 Inference time comparison in Class-IL
@@
L2P Wang et al. (2022e) 9.53
DualPrompt Wang et al. (2022c) 9.51
CODA-P Smith et al. (2023b) 9.47
Adam-ssf Zhou et al. (2024a) 10.03
SEMA Wang et al. (2025) 7.39
@@
C.10 Task-wise resource allocation
@@
Appendix D Experimental setup
D.1 Backbones
To demonstrate the versatility of PEARL across different architectures, we also explore three distinct backbones: (i) Convolutional networks such as ResNet-18 and a ResNet-18-like architecture incorporating Blueprint Separable Convolutions (BSC) (Li et al., 2022b), and (ii) the Vision Transformer (ViT) Base 16 pre-trained on ImageNet-21k.
With regard to the PEARL (R18), the backbone is same as the one used in Mammoth CL benchmark (Buzzega et al., 2020). In line with previous works, the backbone is randomly initialized and trained for CL tasks. However, we introduce PyTorch ModuleLists all throughout to make it easier to add / remove convolutional and BatchNorm layers at will. The ResNet-18 backbone consists of 4 layers with 2 blocks each. Each block in turn consists of two 3x3 convolutional layers with a BatchNorm and ReLU activation in between and after. ResNet-18 also entails shortcut connections in the form of a 3x3 convolutional layer followed by a BatchNorm layer and ReLU activation. Within PEARL (R18), we reserve separate classification and BatchNorm layers for each task. As the backbone is not pre-trained, we treat the first task as the reference task and conduct dynamic allocation from the second task onward. LoRA components are allocated for every convolutional layer. Their rank is dynamically decided after SVD decomposition of respective task vectors. Even after dynamic allocation and further training, LoRA components, BatchNorm layers and a corresponding classification layer are kept separate for each task.
Li et al. (2022b) introduced blueprint separable convolutions (BSC) to optimize and reduce redundancy in convolutional operations. Essentially, BSC entails a 1x1 operation (point-wise filter) followed by a depth-wise convolution layer. This separability reduces the memory and computational footprint resulting far more efficient model. Such a model is well suited for parameter isolation CL approach as it reduces overall model size when dealing with longer task sequences. PEARL (BSC) backbone consists of 3 layers with 2 blocks each. Within each block, the organization of layers is same as that of ResNet-18 except that traditional convolutional layers are replaced by BSC layers. As the BSC backbone is a custom, small network, pre-trained model is not readily available. Therefore, we treat the first task as the reference task and conduct dynamic allocation from the second task onward. LoRA components are allocated for every point-wise filters as they have the majority of the footprint in terms of number of parameters. Since depth-wise filters have relatively lower footprint, we allocate LoRA components and select the rank of the task vector after decomposition. Same as in ResNet-18, even after dynamic allocation and further training, LoRA components, BatchNorm layers and a corresponding classification layer are kept separate for each task.
@@ D.1 Backbones
D.2 Datasets and settings
We evaluate PEARL across two distinct scenarios of CL: Class-IL and Task-IL. In both these settings, the model encounters a new set of classes in each task and must learn to distinguish all classes encountered thus far after each task. In practice, we split CIFAR-100 (Krizhevsky et al., 2009), TinyImageNet (Le & Yang, 2015), and ImageNet-R (Hendrycks et al., 2021) into different tasks depending on the number of tasks as in Table 17.
@@ D.2 Datasets and settings
D.3 Training details
For experiments concerning PEARL (ViT) on the ImageNet-R dataset, we use the Adam optimizer (Kingma & Ba, 2015) with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The batch size is set to 128. The ViT model is fine-tuned on the current task with a learning rate of $5\times 10^{-3}$ for 25 epochs ($E_{1}$). Following this, the LoRA weights are trained separately for an additional 25 epochs ($E_{2}$) using a learning rate of $5\times 10^{-4}$. We observed that a cosine-decaying learning rate schedule outperforms a constant learning rate. Regarding the LoRA parameterization, the rank is determined dynamically based on the criteria outlined in our approach, and the LoRA alpha parameter is set to twice the rank.
With regard to PEARL convolutional variants, we use a batch size of 32 for all experiments. Furthermore, we fine-tune the reference task weights for 15 epochs ($E_{1}$) with a best learning rate found in Section C.3. Following this, the LoRA components are initialized with dynamic ranks and are fine-tuned further for 50 epochs ($E_{2}$).